A particle of mass 10g is placed in potential field given by V= (50x^2+100)erg/g. what will be the frequency of oscillation?

Grade:12th pass

2 Answers

pss
7 years ago
Considering the particle mass of 10g, the actual V = (50x^2 +100 erg)*10 = 500x^2 +1000 erg.
Now the frequency is given by E = h* (freequency)
Freequency =   (500*x^2 + 1000)/6.626 * 10^-27 erg-seconds
Hence Freequency = (75.46*x^2+150.9) * 10^27/sec
RITIK GUPTA
6 years ago
P.E=mV=10(50x^2+100)=500x^2+1000F=-dU/dxF=-1000xF=mama=-1000xa=-1000x/10=-100xa=-w(omega)^2xw=√100=10f=w/2π=10/2π=3.2sec^(-1)